OpenStudy (anonymous):

What is the value of X? 6(-3-x)-2x=14

sammixboo (sammixboo):

First, simplify the equation to get this \(\rm \color{#20bd23}{-8x~-~18~=~14}\) Do you know how to solve that equation?

OpenStudy (anonymous):

no

sammixboo (sammixboo):

Ok, so you first add 18 to both sides of the equation \(\rm \color{#20bd23}{-8x~-~18~+~18=~14~+~18}\) \(\rm \color{#20bd23}{-8x~=~14~+~18}\) What is 14 + 18?

OpenStudy (anonymous):

32

sammixboo (sammixboo):

Correct, so \(\rm \color{#20bd23}{-8x~=~32}\) Now divide both sides of the equation by -8 \(\rm \color{#20bd23}{\dfrac{-8x}{-8}=\dfrac{32}{-8}}\) \(\rm \color{#20bd23}{x=\dfrac{32}{-8}}\) What is 32/-8?